# Dark Matter (H)eats Young Planets

### v2

@@ Abstract
We study the effect of dark matter annihilation on the formation of Jovian planets. We show that dark matter heat injections can slow or halt Kelvin-Helmholtz contraction, preventing the accretion of hydrogen and helium onto the solid core. The existence of Jupiter in our solar system can therefore be used to infer constraints on dark matter with relatively strong interaction cross sections. We derive novel constraints on the cross section for both spin-dependent and spin-independent dark matter. We highlight the possibility of a positive detection using future observations by JWST, which could reveal strongly varying planet morpholoiges close to our Galactic Center.
@@
Our technical capabilities to detect planets outside our solar system have developed dramatically. From the first indirect evidence for an exoplanet Mayor and Queloz (1995), to direct imaging Chauvin et al. (2005) and precision atmospheric observations Martins et al. (2015), we have now 5.5× $5.5\times\,10^{3}$ detected exoplanets, and $\mathcal{O}(10^{4})$ candidates awaiting confirmation cat (2020). Especially with the onset of high performance infra-red astronomy and the JWST Brande et al. (2019), and Roman Johnson et al. (2020) space missions low temperature targets, such as exoplanets can be studied across the galaxy, all the way to the Galactic Center (GC). This development opens up a new window to our galaxy, and allows us to study exoplanets in various astrophysical environments.
@@
Dark Matter (DM) as the dominant mass source in our galaxy is well mapped by stellar kinematics, and despite uncertainties displays a density profile that raises towards the GC Benito et al. (2021). In the search for this ellusive substance a number of studies have considered DM capture in celestial object ranging from the Earth and the Sun Batell et al. (2010); Pospelov et al. (2008); Pospelov and Ritz (2009); Rothstein et al. (2009); Chen et al. (2009); Schuster et al. (2010a, b); Bell and Petraki (2011); Feng et al. (2016a); Kouvaris and Tinyakov (2010a); Feng et al. (2016b); Allahverdi et al. (2017); Leane et al. (2017); Arina et al. (2017); Albert et al. (2018a, b); Nisa et al. (2019); Niblaeus et al. (2019); Cuoco et al. (2020); Serini et al. (2020); Acevedo et al. (2021); Mazziotta et al. (2020); Bell et al. (2021), Jupiter Batell et al. (2010); Leane and Linden (2021); Li and Fan (2022); French and Sher (2022); Ray (2023), Brown Dwarfs Leane and Smirnov (2021); Leane et al. (2021), Uranus Mitra (2004), White Dwarfs Garani et al. (2023), Neutron Stars Goldman and Nussinov (1989); Gould et al. (1990); Kouvaris (2008); Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010b); McDermott et al. (2012); Kouvaris and Tinyakov (2011a); Guver et al. (2014); Bramante et al. (2013); Bell et al. (2013); Bramante et al. (2014); Bertoni et al. (2013); Kouvaris and Tinyakov (2011b); McCullough and Fairbairn (2010); Angeles Perez-Garcia and Silk (2015); Bramante (2015); Graham et al. (2015); Cermeno et al. (2016); Graham et al. (2018); Acevedo and Bramante (2019); Janish et al. (2019); Krall and Reece (2018); Baryakhtar et al. (2017); Raj et al. (2018); Bell et al. (2018); Chen and Lin (2018); Dasgupta et al. (2019); Hamaguchi et al. (2019); Camargo et al. (2019); Bell et al. (2019); Acevedo et al. (2020); Joglekar et al. (2019, 2020); Bell et al. (2020); Garani et al. (2020); Leane et al. (2021); Bramante et al. (2022a); Coffey et al. (2022), and other stars Freese et al. (2009); Taoso et al. (2008); Ilie et al. (2020a, b); Lopes and Lopes (2021); Ellis (2021); Croon and Sakstein (2023), as well as recently Exoplanets Leane and Smirnov (2021); Bramante et al. (2022b) (for a recent review, see Bramante and Raj (2023)). Many of the above studies rely on an energy injection from DM particles that annihilate after being captured by the celestial object. This is expected in models where DM has been thermally produced in the early universe.
@@
Jovian exoplanets have been identified as ideal candidates for DM capture in Ref. Leane and Smirnov (2021). The advantage is their relatively large surface gravity, combined with rather low core temperatures, leading to DM with low masses to be retained in the object. Their hydrogen-rich composition makes them kinematically well suited for capturing DM particles around the GeV scale, and especially particles with spin-dependent nuclear interactions. The differential measurement of Jovian planet and Brown Dwarf temperatures close to the GC has been identified as a promising tool to detect heating from annihilating DM  Leane and Smirnov (2021).
@@
Figure 1: Benchmark for spin-dependent dark matter: $m_{\rm DM}=10$ GeV, $\sigma_{p}=4\times 10^{-28}\,\rm cm^{2}$. The red line shows the dark matter capture fraction at the time when the planet is $10\,\rm M_{\oplus}$ and the black continuous and dashed lines show the fractions needed to halt formation through gas capture. This fraction depends on the temperature of the surrounding gas, and thus on the proximity to the host star.
In Fig. 1 we show the capture efficiency as a function of distance from the GC assuming an NFW profile and a Maxwellian velocity dispersion. As shown here, at a given distance from the host star (corresponding to a given external temperature $T_{\rm ext}$), planet formation can be halted in the GC while it does occur some distance away from it. This would imply that towards the GC, Jovian planets appear to occur further closer to the host stars, modulo migration after formation. This observation provides a new and radically different search for a DM heating signal within our galaxy. This distinct phenomenology can be targeted by future observations of exoplanets leveraging the advances in the development of infrared telescopes.
Finally, the mere existence of Jupiter in our solar system allows us to set new constraints on the parameter space of DM. We find that future observations of exoplanets near the GC, enabled by the JWST and future infrared telescopes will further test yet unconstrained parameter space of DM candidates across a wide range of masses and interaction cross sections.
@@
We focus on giant, Jovian planets, and the formation process called core-accretion gas-capture Hubickyj et al. (2005), which proceeds in two stages. First, a solid core made of heavier elements, such as ice and silica forms within a fraction of a Myr. These planets form beyond the frost line of a solar system, where icy compounds are cool enough to remain solid, allowing cores to become heavy enough to start to accrete hydrogen and helium. Planets formed closer to their sun, such as Earth, on the inside the frost line have cores dominated by iron and silicates, which are too rare to lead to large cores that would accrete gas.
Gas capture does not begin until the core accretion exceeds about $10M_{\oplus}$ ($M_{\oplus}=6.0\times 10^{24}$ kg), when the growing core raises the escape velocity above the thermal velocity of the nebular gas. The proto-planet then accumulates a gas envelope at an increasing rate, reflecting its increasing mass. The gas can quickly fill up the Bondi sphere of the planet, which subsequently needs to cool via the Kelvin-Helmholtz mechanism in order to accrete further Pollack et al. (1996); Adams and Batygin (2022), until eventually the protoplanet reaches a threshold mass at which runaway accretion takes place. Therefore, during the gas accretion period before runaway (which takes place on the scale of a Myr  D’Angelo et al. (2021)) the system remains in a quasi-equilibirum state where gas accretion can only proceed once the excess heat that is released from the gravitational binding energy has been radiated away. It is then not surprising that an additional heat source can disrupt this fragile equilibrium and lead to drastically different outcomes. Note in particular that the hydrogen accretion stage is unique, as hydrogen atoms have the lowest gravitational binding energy and are thus most susceptible to evaporation from the forming planet.
@@
where $\dot{Q}$ is the radiative heat loss, $dU/dt$ the change in gravitational potential energy, $L_{\rm DM}(R)$ is the $R$ dependent heat injection due to DM annihilation, which we call DM luminosity (note that this is not necessarily directly observable as luminosity of the protoplanet), and $1/2\leq c_{1}\leq\infty$ is a factor which depends on the internal mass distribution (for a uniform sphere, $c_{1}=3/5$).
@@
Initially, dark matter heat injections will raise the temperature of the protoplanet, increasing its radiative heat loss. However, above a certain temperature the mixture of hydrogen and helium gas acquires enough thermal energy such that particles near its surface reach the escape energy and become unbound.
This process will be in competition with the mass accretion rate, which we take from numerical simulations Lissauer et al. (2009). To quantify the particle loss rate we consider the generalisation of the Jeans escape rate $\Gamma^{0}_{\rm J}(T)$ Jeans (2009) by including the effect of a chemical potential difference
$\Gamma_{\rm J}(T)=\Gamma^{0}_{\rm J}(T)\exp\left(-\frac{\Delta\mu}{N\,T}\right)$
$=\frac{nv}{2\sqrt{\pi}}\left(1+\frac{v_{\rm esc}^{2}}{v^{2}}\right)\exp\left(-\frac{E_{\rm esc}}{T}\right),$ (2)
where $n$ is the number density of gas particles, $v_{\rm esc}$ the surface escape velocity, $v=\sqrt{2T/m}$, and the escape energy that takes into account the difference in chemical potentials
$E_{\rm esc}=\frac{G_{N}Mm}{r}+\frac{\Delta\mu}{N}=\frac{m\,v_{\rm esc}^{2}}{2}+\frac{\Delta\mu}{N}\,.$ (3)
Here, $G_{N}$ is the Newton constant, $M$ the mass of the forming planet, $m$ the mass of the evaporating particle, and $\Delta\mu/N$ is the difference in chemical potentials per particle. It is intriguing that the chemical potential is relevant in this case, reflecting the special situation of evaporation of particles from the planetary envelope into the surrounding gas of the protoplaneary disk, which has a comparable particle number density and temperature.
Given that $\mu_{1}$ is the chemical potential in the envelope and $\mu_{2}$ that of the protoplaneraty disk, the difference in the chemical potentials $\Delta\mu$ is given by
$\mu_{1}-\mu_{2}=U_{1}-U_{1}+P_{1}V_{1}-P_{2}V_{2}+T_{2}S_{2}-T_{1}S_{1}$ (4)
where we assumed that the gravitational energies of the particles are of the same order. We further assume that the particles behave as ideal gas, and thus the internal energy is $U_{i}=C_{V}N_{i}T_{i}$, $P_{i}V_{i}=N_{i}T_{i}$, and the entropy is given by the Sackur-Tetrode formula $S_{i}=N_{i}\kappa_{i}$, with
$\kappa_{i}=\left(\frac{5}{2}+\log\left[\frac{V_{i}}{N_{i}}\left(\frac{4\pi mU_{i}}{3N_{i}}\right)^{3/2}\right]\right)\,.$ (5)
Furthermore, we assume that the gases are monoatomic, and hence the heat capacity $C_{V}\approx 3/2$.
We can therefore express the chemical potential difference per envelope particle as
$\frac{\Delta\mu}{N_{1}}=C_{V}\left(T_{1}-\frac{N_{2}}{N_{1}}T_{2}\right)+T_{1}-\frac{N_{2}}{N_{1}}T_{2}+\bar{\kappa}\left(T_{2}\frac{N_{2}}{N_{1}}-T_{1}\right)$
$\approx\left(1+C_{V}-\bar{\kappa}\right)\left(T_{1}-T_{2}\right)=\frac{3}{2}\log\left[\frac{4\pi m\bar{u}}{3\bar{n}^{2/3}}\right]\left(T_{2}-T_{1}\right)\,.$ (6)
Where we have used the assumption that $N_{1}\approx N_{2}$, and as the logarithm is a slowly varying function, the argument can be replaced by the average value of the internal energy per particle $\bar{u}=\bar{U}/\bar{N}\approx C_{V}\bar{T}$. Thus, we observe that in the case that the surrounding gas is hotter than the envelope an additional suppression slows particle evaporation down.
Now we compare the escape rate with the rate of growth of the planet in the absence of dark matter heat $\dot{M}$, which we take from simulation results presented in Lissauer et al. (2009). Equating those rates
$A_{\rm planet}\,\Gamma_{\rm J}(T)|_{T=T_{\rm max}}=\mathcal{O}(1)\,\dot{M}$ (7)
where $A_{\rm planet}$ is the surface from which evaporation takes place, we find that the result is insensitive to the value of the numerical $\mathcal{O}(1)$ factor in the equation, since it can not compete with the exponential in the escape rate (see also D’Angelo et al. (2021), which is largely consistent but presented slower accretion, implying smaller DM injections are needed to match it).
For an ambient temperature of $115$ K of the surrounding gas Lissauer et al. (2009), this treatment gives us hydrogen escape temperatures of $T_{\rm max}\sim 85$ K for the envelope. For the core the assumption above has to be modified given that the particle number density in the core is much higher than in the surrounding gas. We get $T_{\rm max}\sim 10^{3}$ K for the core, a result that is not significantly affected by the ambient gas properties for the systems we study.
@@
Thus, the maximum heat loss happens just below $T_{\rm max}$ and we estimate that DM heat injection halts further gas capture if
@@
$\begin{split}L_{\rm DM}&\geq 4\pi R^{2}\sigma T_{\rm max}^{4}\\ &\geq 4\times 10^{-4}\times\left(\frac{R}{10^{3}R_{p}}\right)^{2}\,\left(\frac{T_{\rm max}}{80\rm K}\right)^{4}\rm\,L_{\odot}\end{split}$ (8)
@@
where $R_{p}=7\times 10^{5}$ km is the radius of Jupiter (the protoplenatary envelope has radius $\sim 10^{3}R_{p}$), and where $L_{\odot}=2\times 10^{36}\,\rm GeV\,s^{-1}$ is the luminosity of the Sun. To investigate whether Eq. (8) is satisfied in the evolution of the planet, we use the simulation results presented in D’Angelo et al. (2021). We will now examine how the DM luminosity depends on the properties of the capturing object, and the DM particle model.
@@
$L_{\rm DM}=m_{\rm DM}C_{\rm cap}=m_{\rm DM}f_{\rm cap}\Phi\pi R^{2}$ (9)
where $C_{\rm cap}$ is the capture rate, and $f_{\rm cap}$ the captured fraction of the DM flux $\Phi$ passing through the object, given by
$\Phi=\,v_{\rm DM}\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}}{m_{\rm DM}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\,.$ (10)
@@
$\begin{split}L_{\rm DM}\sim&7\times 10^{-4}f_{\rm cap}\left(\frac{R}{10^{3}R_{p}}\right)^{2}\\ &\times\left(\frac{v_{\rm DM}}{270\rm km\,s^{-1}}\right)\left(\frac{\rho_{\rm DM}}{0.42\,\rm GeV\,cm^{-3}}\right)\,L_{\odot}\end{split}$ (11)
which only depends on the dark matter mass through $f_{\rm cap}$. It is seen that this may indeed exceed (8).
@@
Given a certain DM mass and velocity, the DM particle needs a certain number of scatterings to drop below the escape velocity of a given object. This will in general also depend on the efficiency of the energy loss, determined by the mass ratio of the DM particle and the target nucleus – DM is not sufficiently slowed down when this mass ratio is too large. Furthermore, in the case of DM particles with masses similar to or below the target nucleus mass, DM particles might be reflected off the object before capture can occur, as discussed in detail in Ref. Leane and Smirnov (2023a, b). To compute the fraction of the DM particles that are captured in a given object we use the Asteria Leane and Smirnov (2023b, a) package, which takes all the above mentioned phenomena into account.
@@
where $\mu(m_{\chi},m_{\rm SM})$ is the reduced mass of a DM and a SM particle of mass $m_{\rm SM}$. Note that it has been pointed out that at cross sections $\sigma_{\chi N}>10^{-26}\text{cm}^{2}$ the scattering is not described by interactions of point like particles, and the coherent scattering picture breaks down Digman et al. (2019); Cappiello et al. (2021). Rather, at cross sections larger than this threshold the interaction can be assumed as a total geometric scattering cross section with the entire nucleus without an implied scaling with atomic number.
@@
To compute the capture of DM in the protoplanet we introduce a simplifying effective framework and use the Asteria Leane and Smirnov (2023b, a) package to compute capture in two regimes depending on where in the object we expect the DM to be captured. In the first regime the DM capture takes place on the scale of the entire object and the average chemical composition of the protoplanet is taken as an input for Asteria, while in the other regime the capture takes place in the envelope. In the second case only the chemical composition of the envelope is used, but the total mass of the protoplanet determines the escape velocity. To distinguish the regimes we define the effective capture length $\lambda_{\rm cap}$ and compare it to the size of the protoplanet, which is dominated by the size of the envelope.
$\lambda_{\rm cap}=\lambda_{\rm mfp}\,N_{\rm cap}\text{ with }\lambda_{\rm mfp}=\left(\sigma_{\rm SI/SD}\,n_{\rm SM}\right)^{-1}\,,$ (14)
$N_{\rm cap}=\frac{\log\left(v_{\text{esc}}^{2}/(v_{\chi}^{2}+v_{\text{esc}}^{2})\right)}{\log\left(1-2\mu(1-\mu)^{-2}\right)}\,,$
where $N_{\rm cap}$ is the number of scatterings required to capture a DM particle Mack et al. (2007), $v_{\text{esc}}$ the protoplanets’ escape velocity, $v_{\chi}$ the DM impact velocity, and $\mu=m_{\rm DM}/m_{\rm SM}$ is the mass ratio between the DM and the average SM particle in the protoplanet. Note that this estimate provides an upper bound on the capture length of DM as a straight-line trajectory is assumed.
To determine the chemical composition we model the protoplanet in a simplified manner assuming that it consists of an icy core are an outer gas envelope, both are assumed to have a homogeneous average density. We use the mass and radius from Ref. Lissauer et al. (2009) as time dependent input quantities. The outer envelope has a large radius and a low density. We model the envelope to consist of 75% hydrogen and 25% helium. The core is assumed to have a $58/42$ ice ($\rm H_{2}O$) to silica ($\rm SiO_{2}$) mass ratio D’Angelo et al. (2021), which makes Oxygen the most abundant element. This implies that in the regime where the core contributes to the capture, optimal capture happens for slightly heavier DM candidates, due to more efficient energy transfer between the SM and DM particles of similar mass.
@@
Lastly, we note that light DM candidates can be reflected, such that the capture efficiency remains below one even at large cross sections Leane and Smirnov (2023a). On the other hand, in models with a long range force mediator, the escape velocity in the medium can be increased, which in turn suppressed the reflection probability Acevedo et al. (2023). We discuss the implications for the cross section sensitivity estimate in the following section.
@@
Figure 2: Constraints on spin-dependent parameter space. The filled purple contours demonstrate the cross sections for which the formation of Jupiter stalls at $10\rm\,M_{\oplus}$, before gas is captured (see text for motivation of this benchmark). The green contours demonstrate the cross sections for which the formation of Jovian planets near the GC stalls at this same mass. Continuous contours assume that the environment in which the planet forms is $115\rm K$; dashed contours assume it is $20\rm K$. Reflection of light DM is considered for all but the outer purple $T=115\rm K$ contour. As we discussed above reflection is a model dependent phenomenon, and can be strongly suppressed in models with a long range force mediator. Finally, in the region below the orange dotted line DM evaporates from the forming planet as discussed in appendix A, given short-range interactions only. We also show constraints from direct detection (filled gray contours) and from the CMB Boddy and Gluscevic (2018) and MW-satelites Nadler et al. (2019) (dashed lines).
Figure 3: Constraints on spin-independent DM interactions from the existence of Jupiter, and expected sensitivity from the non-existence of Jovians in the GC, as well as direct detection and cosmology dependent constraints – as in Fig.2.
@@
We study the DM fraction needed to halt the formation process for Jupiter in the solar system, using $\rho_{\rm DM}=0.42\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=270\rm km\,s^{-1}$, and find that $f_{\rm cap}$ remains approximately constant for the period of formation that we are interested in.
In Figs. 2 and 3 we show constraints on the spin-dependent and spin-independent scattering cross sections respectively from the fact that the formation of Jupiter in our solar system has not stopped at a mass of $10\rm M_{\oplus}$ as purple contours. As the temperature increase needed to satisfy (Disrupting Jovian Formation) does grow as a function of time, we expect that formation of Jovian planets in the GC halts at $10\rm M_{\oplus}$ in the parameter range indicated by the contours, and proceeds as normal below these contours. We also project constraints based on the observation of Jovian planets in the GC (for which we have assumed $\rho_{\rm DM}=10^{3}\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=10\rm\,km\,s^{-1}$) in chartreuse (note that the premature ending is already in tension with inferences of Jovian planets in the GC from lensing events Sahu et al. (2006); Janczak et al. (2010); Koshimoto et al. (2021)). We show these constraints and projections for two different benchmarks for the protostellar disk temperature; the more conservative $115$ K and $20$ K, for planets forming further away from the star. We note that this is not the only parameter degeneracy; most notably, the unavailability of nearby gas and dust can also cause a protoplanet to stop accreting, and the migration of a planet away from the star can likewise change its accretion rate. Therefore, while one can interpret the observation of a large gas giant as a constraint on DM, the non-observation cannot directly be interpreted as positive evidence.

### v3

@@ Abstract
We study the effect of dark matter annihilation on the formation of Jovian planets. We show that dark matter heat injections can slow or halt Kelvin-Helmholtz contraction, preventing the accretion of hydrogen and helium onto the solid core. The existence of Jupiter in our solar system can therefore be used to infer constraints on dark matter with relatively strong interaction cross sections. In the case of spin-dependent dark matter, we derive novel constraints beyond the reach of current direct detection experiments. We highlight the possibility of a positive detection using future observations by JWST, which could reveal strongly varying planet morpholoiges close to our Galactic Center.
@@
Our technical capabilities to detect planets outside our solar system have developed dramatically. From the first indirect evidence for an exoplanet 1995Natur.378..355M, to direct imaging 2005A&A...438L..29C and precision atmospheric observations 2015A&A...576A.134M, we have now 5.5× $5.5\times\,10^{3}$ detected exoplanets, and $\mathcal{O}(10^{4})$ candidates awaiting confirmation catalog. With the onset of high performance infra-red astronomy and the JWST Brande_2019, and Roman Johnson_2020 space missions low temperature targets, such as exoplanets can be studied across the galaxy, all the way to the Galactic Center (GC). This development opens up a new window to our galaxy, and allows us to study exoplanets in various astrophysical environments.
@@
Dark Matter (DM) as the dominant mass source in our galaxy is well mapped by stellar kinematics, and despite uncertainties displays a density profile that raises towards the GC Benito:2020lgu. In the search for this ellusive substance a number of studies have considered DM capture in celestial object ranging from solar system objects such as the Earth, Sun Batell:2009zp; Pospelov:2007mp; Pospelov:2008jd; Rothstein:2009pm; Chen:2009ab; Schuster:2009au; Schuster:2009fc; Bell_2011; Feng:2015hja; Kouvaris:2010; Feng:2016ijc; Allahverdi:2016fvl; Leane:2017vag; Arina:2017sng; Albert:2018jwh; Albert:2018vcq; Nisa:2019mpb; Niblaeus:2019gjk; Cuoco:2019mlb; Serini:2020yhb; Acevedo:2020gro; Mazziotta:2020foa; Bell:2021pyy, Jupiter Batell:2009zp; Leane:2021tjj; Li:2022wix; French:2022ccb; Ray:2023auh, and Uranus Mitra:2004fh, to Brown Dwarfs Leane:2020wob; Leane:2021ihh, White Dwarfs Garani:2023esk, Neutron Stars Goldman:1989nd; Gould:1989gw; Kouvaris:2007ay; Bertone:2007ae; deLavallaz:2010wp; Kouvaris:2010vv; McDermott:2011jp; Kouvaris:2011fi; Guver:2012ba; Bramante:2013hn; Bell:2013xk; Bramante:2013nma; Bertoni:2013bsa; Kouvaris:2010jy; McCullough:2010ai; Perez-Garcia:2014dra; Bramante:2015cua; Graham:2015apa; Cermeno:2016olb; Graham:2018efk; Acevedo:2019gre; Janish:2019nkk; Krall:2017xij; Baryakhtar:2017dbj; Raj:2017wrv; Bell:2018pkk; Chen:2018ohx; Dasgupta:2019juq; Hamaguchi:2019oev; Camargo:2019wou; Bell:2019pyc; Acevedo:2019agu; Joglekar:2019vzy; Joglekar:2020liw; Bell:2020jou; Garani:2020wge; Leane:2021ihh; Bramante:2021dyx; Coffey:2022eav; Nguyen:2022zwb; Alvarez:2023fjj, and other stars Freese:2008hb; Taoso:2008kw; Ilie:2020iup; Ilie:2020nzp; Lopes:2021jcy; Ellis:2021ztw; DCandJS, as well as recently Exoplanets Leane:2020wob; Bramante:2022pmn. For a recent review of searches in compact stars, see Bramante:2023djs. Many of the above studies rely on an energy injection from DM particles that annihilate after being captured by the celestial object. This is expected in models where DM has been thermally produced in the early universe.
@@
Jovian exoplanets have been identified as ideal candidates for DM capture in Ref. Leane:2020wob. The advantage is their relatively large surface gravity, combined with rather low core temperatures, leading to DM with low masses to be retained in the object. Their hydrogen-rich composition makes them kinematically well suited for capturing DM particles around the GeV scale, and especially particles with spin-dependent nuclear interactions. The differential measurement of Jovian planet and Brown Dwarf temperatures close to the GC has been identified as a promising tool to detect heating from annihilating DM  Leane:2020wob.
@@
Figure 1: Captured DM fraction in a benchmark scenario for spin-dependent dark matter: $m_{\rm DM}=5$ GeV, $\sigma_{p}=10^{-27}\,\rm cm^{2}$. Dark matter density and velocity as in the solar system. The red line gives the fraction of the dark matter flux through the planet that is captured by the core, the cyan line gives the fraction captured by the envelope, and the dashed line the total captured fraction. Above the black line labeled $T>T_{\rm max}$ dark matter annihilation heats the planet above the temperature defined by Eq. (2), which halts the formation process.
The mere existence of Jupiter in our solar system allows us to set new constraints on the parameter space of DM, which are especially competitive in the case of spin-dependent interactions. We find that future observations of exoplanets near the GC, enabled by the JWST and future infrared telescopes will further test yet unconstrained parameter space of DM candidates across a wide range of masses and interaction cross sections.
@@
We focus on giant, Jovian planets, and the formation process called core-accretion gas-capture 2005Icar..179..415H, which proceeds in two stages. First, a solid core made of heavier elements, such as ice and silica forms within a fraction of a Myr. These planets form beyond the frost line of a solar system, where icy compounds are cool enough to remain solid, allowing cores to become heavy enough to start to accrete hydrogen and helium. Planets formed closer to their sun, such as Earth, on the inside the frost line have cores dominated by iron and silicates, which are too rare to lead to large cores that would accrete gas.
Gas capture does not begin until the core accretion exceeds about $10M_{\oplus}$ ($M_{\oplus}=6.0\times 10^{24}$ kg), when the growing core raises the escape velocity above the thermal velocity of the nebular gas. The proto-planet then accumulates a gas envelope at an increasing rate, reflecting its increasing mass. The gas can quickly fill up the Bondi sphere of the planet, which subsequently needs to cool via the Kelvin-Helmholtz mechanism in order to accrete further 2022DPS....5410202A. Therefore, during the gas accretion period (which takes place on the scale of a Myr  2021Icar..35514087D) the system remains in a quasi-equilibirum state where gas accretion can only proceed once the excess heat that is released from the gravitational binding energy has been radiated away. It is then not surprising that an additional heat source can disrupt this fragile equilibrium and lead to drastically different outcomes. Note in particular that the hydrogen accretion stage is unique, as hydrogen atoms have the lowest gravitational binding energy and are thus most susceptible to evaporation from the forming planet.
@@
where $\dot{Q}$ is the radiative heat loss, $dU/dt$ the change in gravitational potential energy, $L_{\rm DM}(R)$ is the $R$ dependent heat injection due to DM annihilation, which we call DM luminosity, and $1/2\leq c_{1}\leq\infty$ is a factor which depends on the internal mass distribution (for a uniform sphere, $c_{1}=3/5$).
@@
Initially, dark matter heat injections will raise the temperature of the protoplanet, increasing its radiative heat loss. However, above a certain temperature the mixture of hydrogen and helium gas acquires enough thermal energy such that particles near its surface reach the escape velocity. We estimate that this happens for temperatures above which
$\Gamma_{\rm J}(T)=\left.\frac{nv}{2\sqrt{\pi}}\left(1+\frac{v_{\rm esc}^{2}}{v^{2}}\right)\exp\left(-\frac{v_{\rm esc}^{2}}{v^{2}}\right)\right|_{T=T_{\rm max}}=C\dot{M}$ (2)
where $\Gamma_{\rm J}$ is the flux of escaping particles given by the Jeans escape formula jeans_2009, with $n$ the number density of gas particles, $v_{\rm esc}$ the surface escape velocity, $v=\sqrt{2T/m}$, with $m$ being the mass of the gas particle, and $\dot{M}$ is the rate of growth of the planet in the absence of dark matter heat, which we take from simulation results presented in Lissauer:2008hn; d2021growth. Equation (2) overestimates the actual loss rate because it assumes the tail of the velocity distribution is replenished instantaneously, Therefore, to be conservative, we take $C=10$. This gives us maximum hydrogen escape temperatures of $T_{\rm max}\sim 8$ K and $10^{3}$ K in the envelope and in the core respectively.
@@
Thus, the maximum heat loss happens just below $T_{\rm max}$ and we estimate that DM heat injection halts collapse if
@@
$\begin{split}L_{\rm DM}&\geq 4\pi R^{2}\sigma T_{\rm max}^{4}\\ &\geq 4\times 10^{-8}\times\left(\frac{R}{10^{3}R_{p}}\right)^{2}\,\left(\frac{T_{\rm max}}{8\rm K}\right)^{4}\rm\,L_{\odot}\end{split}$ (3)
@@
where $R_{p}=7\times 10^{5}$ km is the radius of Jupiter (the protoplenatary envelope has radius $\sim 10^{3}R_{p}$), and where $L_{\odot}=2\times 10^{36}\,\rm GeV\,s^{-1}$ is the luminosity of the Sun. To investigate whether Eq. (3) is satisfied in the evolution of the planet, we use the simulation results presented in d2021growth. We will now examine how the DM luminosity depends on the properties of the capturing object, and the DM particle model.
@@
$L_{\rm DM}=m_{\chi}C_{\rm cap}=m_{\chi}f_{\rm cap}\Phi$ (4)
where $m_{\chi}$ is the DM mass, $C_{\rm cap}$ is the capture rate, and $f_{\rm cap}$ the captured fraction of the DM flux $\Phi$ passing through the object, given by
$\Phi=v_{\rm DM}\sqrt{\frac{8}{3\pi}}\frac{\rho_{\rm DM}}{m_{\chi}}\left(1+\frac{3}{2}\frac{v_{\rm esc}^{2}}{v_{\rm DM}^{2}}\right)\,.$ (5)
@@
$\begin{split}L_{\rm DM}\sim&2\times 10^{-9}f_{\rm cap}\left(\frac{R}{10^{3}R_{p}}\right)\left(\frac{M}{10M_{\oplus}}\right)\\ &\times\left(\frac{v_{\rm DM}}{270\rm km\,s^{-1}}\right)^{-1}\left(\frac{\rho_{\rm DM}}{0.42\,\rm GeV\,cm^{-3}}\right)\,L_{\odot}\end{split}$ (6)
independently of the dark matter mass. It is seen that this may exceed (3) for greater $M$ or $\rho_{\rm DM}$.
@@
Given a certain DM mass and velocity, the DM particle needs a certain number of scatterings to drop below the escape velocity of a given object. This will in general also depend on the efficiency of the energy loss, determined by the mass ratio of the DM particle and the target nucleus. Furthermore, in the case of DM particles with masses similar of below the target mass, DM particles might be reflected off the object before capture can occur, as discussed in detail in Ref. RKLandJS; asteria. To compute the fraction of the DM particles that are captured in a given object we use the Asteria asteria; RKLandJS package, which takes all the above mentioned phenomena into account.
@@
where $\mu(m_{\chi},m_{\rm SM})$ is the reduced mass of a DM and a SM particle of mass $m_{\rm SM}$. Note that it has been pointed out that at cross sections $\sigma_{\chi N}>10^{-26}\text{cm}^{2}$ the scattering is not described by interactions of point like particles, and the coherent scattering picture breaks down Digman:2019wdm; Cappiello:2020lbk. Rather, at cross sections larger than this threshold the interaction can be assumed as a total geometric scattering cross section with the entire nucleus without an implied scaling with atomic number.
@@
We model the protoplanet in a simplified manner assuming that it consists of an icy core are an outer gas envelope, both are assumed to have a homogeneous average density. We use the mass and radius from Ref. Lissauer:2008hn as time dependent input quantities. The outer envelope has a large radius and a low density. We model the envelope to consist of 75% hydrogen and 25% helium and compute the fraction of dark matter that is captured. We find that only at substantial cross sections above $10^{-28}\text{cm}^{2}$ a substantial fraction of the DM particles are captured. The most efficient capture rate takes place in the GeV mass range, when the DM mass matches the hydrogen target mass. The core of the protoplanet is much more dense, which has the potential to increase the DM capture rate. We model the core as having a $58/42$ ice ($\rm H_{2}O$) to silica ($\rm SiO_{2}$) mass ratio 2021Icar..35514087D, which makes Oxygen the most abundant element. This implies that optimal capture happens for slightly heavier DM candidates, due to more efficient energy transfer between the SM and DM particles of similar mass.
@@
Lastly, we note that light DM candidates can be reflected, such that the capture efficiency remains below one even at large cross sections RKLandJS. As we will see below, this has implications for the sensitivity to subfractions of DM.
@@
Figure 2: Constraints on spin-dependent parameter space. The filled contour demonstrates the cross sections for which the formation of Jupiter stalls at $30\rm\,M_{\oplus}$. The dashed contours demonstrate the cross sections for which the formation of Jovian planets near the GC stalls at $10\rm\,M_{\oplus}$ (see text for motivation of these conditions). The reflection of light DM candidates implies that the capture efficiency asymptotes. Hence, for a subfraction $f_{\rm DM}=10^{-3}$ (shown in chartreuse) there are no constraints below $m_{\chi}\sim 100$ MeV (the asymptotic value is below the value needed to realise (3)). Finally, in the region below the orange dotted line DM evaporates from the forming planet as discussed in appendix A, given short-range interactions only.
Figure 3: Constraints on spin-independent DM interactions from the existence of Jupiter, and expected sensitivity from the non-existence of Jovians in the GC, as well as current direct detection constraints– as in Fig.2.
@@
In Fig. 1 we show the DM fraction needed to halt the formation process for Jupiter in the solar system, using $\rho_{\rm DM}=0.42\,\rm GeV\,cm^{-3}$ and $v_{\rm DM}=270\rm km\,s^{-1}$, we find that $f_{\rm cap}$ increases faster than the fraction necessary to satisfy (3) as a function of time. This does not go on indefinitely, because the rate of gas accumulation rises with increasing mass. Formation studies have shown that after the planet has grown to about $30M_{\oplus}$ gas capture rises exponentially, however as is seen for the benchmark in Fig. 1 the annihilation of DM captured by the envelope heats the protoplanet enough to halt the formation at about $1.1\,\rm Myr$, when it is approximately $20\rm M_{\oplus}$.